OpenStudy (anonymous):

integrate 4/(x-6)^3 from 6 to 8

OpenStudy (anonymous):

it is improper integral night i see

OpenStudy (anonymous):

yes it is!

OpenStudy (anonymous):

this one doesn't converge either

OpenStudy (anonymous):

find the anti derivative, it is \[\frac{-2}{(x-6)^2}\]

OpenStudy (anonymous):

now replacing x by 8 is no problem, you just get a number, namely \[-\frac{1}{2}\]

OpenStudy (anonymous):

but if you try to take the limit \[\lim_{r\rightarrow 6^+}-\frac{2}{(r-6)^2}\] you get minus infinity, so integral does not exist